OpenStudy (anonymous):

Please help. Find the domain of the following function: g(x)=5x/x^3+16x (answer using interval notation). Can't remember how to do this, help would be greatly appreciated! :)

OpenStudy (anonymous):

Set the denominator to 0 and solve for x. The domain will be all numbers except those values. \[x^{3}+16x=0\] \[x(x ^{2}+16)=0\] x² + 16 = 0 has no real solution, so the only value that makes the denominator be 0 is x = 0. So in interval notation the domain is \[\left( -\infty,0 \right)\cup \left( 0, \right)\]

OpenStudy (anonymous):

awesome. thank you so much!!! =D